a man sells a watch for 75 Rs and gain as much percent as the watch cost him . find the cost price of watch.
 

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
Let cost of price = x
 
Then
 
75-x/x * 100 = x
 
7500 - 100x = x²
 
x² + 100x - 7500 = 0
 
Hence 
 
x = -150 or x = 50
 
Hence x = 50 is correct